two equal positive charges are kept at points A and B.the electric potential at that points between A and B (excluding these points) is studied while moving from A to B.the potential-
  1. continuously increases 
  2. continuously decreases
  3. increases then decreases
  4. decreases then increases

Grade:12

1 Answers

Arun
4 years ago
The potential due to a charge Q is kQ/r, where r is the distance of the point from the charge and k is the constant 1/(4πε).
Let the distance between A and B be dd. So the potential (V) of a point at a distance rr along the line joining A and B is (kq/r)+(kq/(d−r)) , where q is the charge of A and B.
 
Therefore,
V(r)=kq(1/r+1/(d−r))=kqd/(r.(d−r))
 
The graph of V with r is - parabola (y=kx²)
 
Where the potential is along the y axis and the distance r is along the x axis.
 
So we can see that the potential decreases and then increases on moving from A to B.